OpenStudy (anonymous):

Let F(x)=f(g(x)) , where f(−8)=7,f ′ (−8)=4,f ′ (5)=10,g(5)=−8 , and g ′ (5)=−7 , find F ′ (5)

OpenStudy (anonymous):

\[F'(5)=f'(g(5))\times g'(5)\]

OpenStudy (anonymous):

plug in the numbers, see what you get \[f'(g(5))g'(5)=f'(-8)\times(-7)=4\times -7=-28\]